# Quasinormal spectra of a wormhole family: overtone features and a parameter-controlled redshift

### v1

@@ VI.3 Bronnikov–Ellis geometry with asymptotic mass:
VI.4 Bronnikov–Ellis geometry with a chosen redshift function:
@@
with $a_{0}=\frac{2b_{0}}{b_{0}+2M}$, $a_{1}=\frac{4M}{b_{0}+2M}$, $k_{0}=sin^{-1}\sqrt{\frac{\left(b_{0}+2M\right)\left(r-b_{0}\right)}{2b_{0}(r-2M)}}$, $k_{1}=\sqrt{\frac{r\left(b_{0}-2M\right)\left(b_{0}^{2}-4M^{2}\right)\left(r^{2}-b_{0}^{2}\right)}{b_{0}^{3}(r-2M)^{3}}}$, $E(...)=EllipticE$, $F(...)=EllipticF$ and $\Pi(...)=EllipticPi$ function.
@@
As shown in the figures below (Figs. 45, 45), the effective potential changes from being a single barrier to a double barrier for higher $M$. The potentials have been plotted numerically.
@@
For this geometry, as we increase the value of $M$, both the real and imaginary parts of the QNM spectra decrease (See Figs. 47 ,47). Thus, an increase in $M$ causes a redshift. On careful observation, one notices that for lower values of $M$, the real parts of quasinormal modes decrease with the overtone number. However, this pattern reverses when we go to higher $M$ values. The height of the effective potential decreases as we keep on increasing $M$, and after a certain value, the effective potential becomes double-peaked. This may explain the above feature of the quasi-normal spectra.
@@
Similar to the scalar perturbation, the effective potential due to axial perturbation changes from being a single barrier to a double barrier for higher $M$ (See Figs. 49 ,49).
@@
The quasi-normal spectra for axial perturbation have a similar redshift effect as the QNMs for scalar perturbation (Figs. 51 ,51). However, unlike the scalar perturbation case, here, a linear fit does not estimate the $\omega_{QNM}$ of the same overtone order.
For this modified Bronnikov-Ellis wormhole, the spectra change with increasing $M$. For the $M=0$ case, the real part decreases with increasing overtone number, but this trend gradually reverses with increasing $M$. Though both $Re(\omega)$ and $|Im(\omega)|$ decrease with $M$ (similar effect as increasing $b_{0}$ of the original Bronnikov-Ellis spacetime), the detection of the first few overtones can distinguish this geometry (with non-zero redshift function) from the original Bronnikov-Ellis wormhole.
@@
3 0 $2.95241-0.40998i$ $2.65202-0.34116i$ $2.31704-0.25890i$ $1.93583-0.15765i$ $1.47895-0.04526i$
1 $2.86643-1.25767i$ $2.59943-1.05329i$ $2.32019-0.81701i$ $2.02320-0.54789i$ $1.68143-0.25349i$
2 $2.66056-2.19085i$ $2.47232-1.84041i$ $2.28366-1.46458i$ $2.09262-1.06263i$ $1.83772-0.62125i$
5 0 $5.15482-0.46725i$ $4.61743-0.39009i$ $4.01226-0.29556i$ $3.31317-0.17059i$ $2.47482-0.02713i$
1 $5.07286-1.40857i$ $4.56144-1.17693i$ $4.00038-0.89951i$ $3.38786-0.56119i$ $2.69934-0.18747i$
2 $4.90746-2.37118i$ $4.44773-1.98388i$ $3.96646-1.53674i$ $3.46529-1.03582i$ $2.86794-0.49395i$
3 $4.65661-3.37224i$ $4.27300-2.82560i$ $3.89988-2.21918i$ $3.51728-1.57299i$ $3.03059-0.87000i$
7 0 $7.24818-0.48308i$ $6.48787-0.40372i$ $5.62868-0.30567i$ $4.62788-0.17089i$ $3.42055-0.01444i$
1 $7.18438-1.45261i$ $6.44308-1.21430i$ $5.61809-0.92386i$ $4.69804-0.55209i$ $3.66405-0.13378i$
2 $7.05634-2.43256i$ $6.35314-2.03436i$ $5.59222-1.56056i$ $4.78041-1.00184i$ $3.84139-0.39161i$
3 $6.86588-3.43084i$ $6.21733-2.87112i$ $5.54494-2.22360i$ $4.84772-1.50282i$ $4.01693-0.72270i$
9 0 $9.30171-0.48964i$ $8.32354-0.40939i$ $7.21638-0.30982i$ $5.92045-0.16947i$ $4.34739-0.00712i$
1 $9.25035-1.47100i$ $8.28712-1.23004i$ $7.20768-0.93382i$ $5.98667-0.54158i$ $4.60839-0.09362i$
2 $9.14741-2.45870i$ $8.21414-2.05635i$ $7.18761-1.56991i$ $6.07049-0.97279i$ $4.79271-0.31009i$
3 $8.99239-3.45726i$ $8.10424-2.89222i$ $7.15227-2.22372i$ $6.14625-1.44932i$ $4.97551-0.60743i$
Table 7: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$. [The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al. (2024). We extended this to aial perturbation and double-peaked potentials (Higher $M$)]. [Results obtained using the spectral method ($b_{0}=1$).]
VII Remarks and future directions
@@
$\bullet$ We extend the ultrastatic, generalised (two parameter–$b_{0}$, $n>2$) Bronnikov-Ellis wormhole formulated in Kar et al. (1995) by constructing a modified metric with a non-zero asymptotic mass $M$ (furthering the analysis in Blázquez-Salcedo et al. (2018b) for $n=2$). The geometry, energy-momentum tensor and the energy conditions for the new three-parameter ($b_{0}$, $n$, $M$) space-times are discussed briefly, followed by the computation of the corresponding effective potentials for scalar and axial gravitational perturbations.
@@
$\bullet$ Finally, we focused on the $n=2$ wormhole family with asymptotic mass $M$ and observed how the increase in $M$ causes a systematic redshift (lowering of values of $Re(\omega)$). We further observed this redshift effect (w.r.t. changing $M$) for the original Bronnikov–Ellis wormhole but with a chosen redshift function ($e^{2\Psi(r)}=1-\frac{2M}{r}$) Alfaro et al. (2024). The distinguishability of these geometries from their massless counterparts (with larger $b_{0}$) has also been mentioned.
@@ ACKNOWLEDGEMENTS
Appendix A QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
@@
In this ultra-static spacetime, if $(\omega_{R}-i\omega_{I})$ is a solution, then $(-\omega_{R}-i\omega_{I})$ is also a solution, both having the same amplitude. In the tables of the following sections, we show only the solution with a positive real part and double its amplitude (accounting for the amplitude of both solutions).
@@
0 0 $0.68137-0.61775i$ $0.68402-0.61594i^{*}$- -- -- --
1 $0.49683-2.17988i$- -- -- -- --
1 0 $1.57271-0.52970i$ $1.57274-0.52969i^{*}$- -- -- --
1 $1.25957-1.70681i$- -- -- -- --
2 0 $2.54666-0.51267i$ $2.54925-0.51285i$ $0.35745915$ $1.73769-0.30514i$ $1.73771-0.30513i^{*}$ $0.2689666$
1 $2.35300-1.57522i$- -- $1.71239-1.02010i$- --
2 $1.94641-2.75472i$- -- $1.52495-2.03050i$- --
3 0 $3.53429-0.50692i$ $3.53443-0.50686i$ $0.30122677$ $2.95241-0.40998i$ $2.95253-0.41006i$ $0.29811729$
1 $3.39038-1.53568i$- -- $2.86643-1.25767i$- --
2 $3.09988-2.62324i$ $3.09682-2.62481i$ $0.16837004$ $2.66056-2.19085i$ $2.67518-2.17517i$ $0.171151$
4 0 $4.52705-0.50433i$ $4.52734-0.50424i$ $0.26517902$ $4.07625-0.44911i$ $4.07646-0.44904i$ $0.2640608$
1 $4.41536-1.52340i$- -- $3.99432-1.36285i$- --
2 $4.19023-2.57208i$ $4.19075-2.57242i$ $0.1397532$ $3.80100-2.30237i$ $3.79816-2.30549i$ $0.14061934$
5 0 $5.52231-0.50296i$ $5.52282-0.50282i$ $0.23964086$ $5.15482-0.46725i$ $5.15524-0.46714i$ $0.239214$
1 $5.42996-1.51789i$ $5.41796-1.52258i$ $2.009\times 10^{-5}$ $5.07306-1.40721i$- --
2 $5.24705-2.54683i$ $5.25123-2.54327i$ $0.12190640$ $4.90770-2.37002i$ $4.90815-2.37157i$ $0.12362226$
6 0 $6.51897-0.50214i$ $6.51980-0.50195i$ $0.22033499$ $6.20882-0.47711i$ $6.20954-0.47695i$ $0.2201443$
1 $6.44246-1.50964i$ $6.44313-1.53221i$ $1.886\times 10^{-5}$ $6.13729-1.43537i$- --
2 $6.28646-2.53346i$ $6.28819-2.53269i$ $0.11151401$ $5.99151-2.40927i$ $5.99267-2.40806i$ $0.1115693$
7 0 $7.51649-0.50162i$ $7.51934-0.50106i$ $0.20454891$ $7.24818-0.48307i$ $7.24931-0.48285i$ $0.20497677$
1 $7.44933-1.50895i$ $7.45576-1.50817i$ $2.309\times 10^{-5}$ $7.24818-0.48307i$- --
2 $7.31541-2.52544i$ $7.31985-2.52412i$ $0.10156691$ $7.05635-2.43256i$ $7.05785-2.43166i$ $0.10294975$
Table 8: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=2$) under scalar perturbations for $m=0$ $7$ and axial perturbations for $m=2$ $7$. [The QNM spectra for the $n=2$ case were published in Kim (2008); Batic and Dutykh (2024); Blázquez-Salcedo et al. (2018b); Alfaro et al. (2024)]

### v2

@@ V.1 Bronnikov–Ellis geometry with asymptotic mass:
V.2 Bronnikov–Ellis geometry with a chosen redshift function:
@@
with $a_{0}=\frac{2b_{0}}{b_{0}+2M}$, $a_{1}=\frac{4M}{b_{0}+2M}$, $k_{0}=sin^{-1}\sqrt{\frac{\left(b_{0}+2M\right)\left(r-b_{0}\right)}{2b_{0}(r-2M)}}$, $k_{1}=\sqrt{\frac{r\left(b_{0}-2M\right)\left(b_{0}^{2}-4M^{2}\right)\left(r^{2}-b_{0}^{2}\right)}{b_{0}^{3}(r-2M)^{3}}}$, E()=EllipticEE(...)=EllipticE, F()=EllipticFF(...)=EllipticF and $\Pi(...)=EllipticPi$ function.
@@
As shown in the figures below (Figs. 41, 41), the effective potential changes from being a single barrier to a double barrier for higher $M$. The potentials have been plotted numerically.
@@
For this geometry, as we increase the value of $M$, both the real and imaginary parts of the QNM spectra decrease (See Figs. 43 ,43). Thus, an increase in $M$ causes a redshift. On careful observation, one notices that for lower values of $M$, the real parts of quasinormal modes decrease with the overtone number. However, this pattern reverses when we go to higher $M$ values. The height of the effective potential decreases as we keep on increasing $M$, and after a certain value, the effective potential becomes double-peaked. This may explain the above feature of the quasi-normal spectra.
@@
Similar to the scalar perturbation, the effective potential due to axial perturbation changes from being a single barrier to a double barrier for higher $M$ (See Figs. 45 ,45).
The quasi-normal spectra for axial perturbation have a similar redshift effect as the QNMs for scalar perturbation (Figs. 47 ,47). However, unlike the scalar perturbation case, here, a linear fit does not estimate the $\omega_{QNM}$ of the same overtone order.
@@
3 0 2.952410.40998i2.95241-0.40998i 2.652020.34116i2.65202-0.34116i 2.317040.25890i2.31704-0.25890i 1.935830.15765i1.93583-0.15765i 1.478950.04526i1.47895-0.04526i
1 2.866431.25767i2.86643-1.25767i 2.599431.05329i2.59943-1.05329i 2.320190.81701i2.32019-0.81701i 2.023200.54789i2.02320-0.54789i 1.681430.25349i1.68143-0.25349i
2 2.660562.19085i2.66056-2.19085i 2.472321.84041i2.47232-1.84041i 2.283661.46458i2.28366-1.46458i 2.092621.06263i2.09262-1.06263i 1.837720.62125i1.83772-0.62125i
5 0 5.154820.46725i5.15482-0.46725i 4.617430.39009i4.61743-0.39009i 4.012260.29556i4.01226-0.29556i 3.313170.17059i3.31317-0.17059i 2.474820.02713i2.47482-0.02713i
1 5.072861.40857i5.07286-1.40857i 4.561441.17693i4.56144-1.17693i 4.000380.89951i4.00038-0.89951i 3.387860.56119i3.38786-0.56119i 2.699340.18747i2.69934-0.18747i
2 4.907462.37118i4.90746-2.37118i 4.447731.98388i4.44773-1.98388i 3.966461.53674i3.96646-1.53674i 3.465291.03582i3.46529-1.03582i 2.867940.49395i2.86794-0.49395i
3 4.656613.37224i4.65661-3.37224i 4.273002.82560i4.27300-2.82560i 3.899882.21918i3.89988-2.21918i 3.517281.57299i3.51728-1.57299i 3.030590.87000i3.03059-0.87000i
7 0 7.248180.48308i7.24818-0.48308i 6.487870.40372i6.48787-0.40372i 5.628680.30567i5.62868-0.30567i 4.627880.17089i4.62788-0.17089i 3.420550.01444i3.42055-0.01444i
1 7.184381.45261i7.18438-1.45261i 6.443081.21430i6.44308-1.21430i 5.618090.92386i5.61809-0.92386i 4.698040.55209i4.69804-0.55209i 3.664050.13378i3.66405-0.13378i
2 7.056342.43256i7.05634-2.43256i 6.353142.03436i6.35314-2.03436i 5.592221.56056i5.59222-1.56056i 4.780411.00184i4.78041-1.00184i 3.841390.39161i3.84139-0.39161i
3 6.865883.43084i6.86588-3.43084i 6.217332.87112i6.21733-2.87112i 5.544942.22360i5.54494-2.22360i 4.847721.50282i4.84772-1.50282i 4.016930.72270i4.01693-0.72270i
9 0 9.301710.48964i9.30171-0.48964i 8.323540.40939i8.32354-0.40939i 7.216380.30982i7.21638-0.30982i 5.920450.16947i5.92045-0.16947i 4.347390.00712i4.34739-0.00712i
1 9.250351.47100i9.25035-1.47100i 8.287121.23004i8.28712-1.23004i 7.207680.93382i7.20768-0.93382i 5.986670.54158i5.98667-0.54158i 4.608390.09362i4.60839-0.09362i
2 9.147412.45870i9.14741-2.45870i 8.214142.05635i8.21414-2.05635i 7.187611.56991i7.18761-1.56991i 6.070490.97279i6.07049-0.97279i 4.792710.31009i4.79271-0.31009i
3 8.992393.45726i8.99239-3.45726i 8.104242.89222i8.10424-2.89222i 7.152272.22372i7.15227-2.22372i 6.146251.44932i6.14625-1.44932i 4.975510.60743i4.97551-0.60743i
Table 7: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.1b_{0}$, $0.2b_{0}$, $0.3b_{0}$ and $0.4b_{0}$. [The QNM spectra for this geometry (single-peaked scalar potentials) were published in Alfaro et al. (2024). We extended this to axial perturbation and double-peaked potentials (Higher $M$)]. [Results obtained using the spectral method ($b_{0}=1$).]
[…]
VI Remarks and future directions
@@
[…]
@@
[…]
@@ ACKNOWLEDGEMENTS
Appendix A Methods used to find the QNMs
A.1 Time domain and Prony fit
The time domain profile represents the temporal evolution of the field $\psi$ in a fixed background geometry. It was originally formulated in Gundlach et al. (1994). In section II.4, we derived the master wave equation, which may also be rewritten without the separability ansatz for the field. Thus, the equation, using the new coordinates $u$ and $v$, such that $du=dt-dx,\hskip 5.69046ptdv=dt+dx$, reads Konoplya and Zhidenko (2011):
$4\frac{\partial^{2}\psi}{\partial u\partial v}+V(u,v)\psi=0.$ (22)
where $S=(u,v)$, $W=(u+h,v)$, $E=(u,v+h)$, and $N=(u+h,v+h)$ and $h$ is the integration step size for both $u$ and $v$. Eqn. (23) enables us to compute $\psi$ iteratively on the uvu-v grid based on the initial data specified on null surfaces $u=u_{0}$ and $v=v_{0}$. The values of $\psi(N)$, thus calculated, give us the time-domain profile for the given potential.
The time-domain profile of $\psi$ is then fitted using the Prony method $\psi(t)\approx\sum_{j=1}^{p}C_{j}e^{-i\omega_{j}t}$. Here $\omega_{j}$ are the quasinormal frequencies and $C_{j}$ are the corresponding amplitudes. We assume that the ringdown starts at $t=0$ and ends at $t=N_{1}h$, and approximate the profile data as a sum of $p\hskip 2.84544pt(\leq\frac{N_{1}+1}{2})$ terms in the following way:
$x_{n}\approx\psi(nh)=\sum_{j=1}^{p}C_{j}z_{j}^{n},\hskip 56.9055ptz_{j}=e^{-i\omega_{j}h}$ (24)
where $a_{0}=1$. This leads to the matrix equation:
$\begin{bmatrix}x_{p-1}&x_{p-2}&\ldots&x_{0}\\ x_{p}&x_{p-1}&\ldots&x_{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N_{1}-1}&x_{N_{1}-2}&\ldots&x_{N_{1}-p}\end{bmatrix}\begin{bmatrix}a_{1}\\ a_{2}\\ \vdots\\ a_{p}\end{bmatrix}=-\begin{bmatrix}x_{p}\\ x_{p+1}\\ \vdots\\ x_{N_{1}}\end{bmatrix}\hskip 28.45274pt\implies X\cdot a=-x,$ (26)
$\omega_{j}=i\frac{\ln(z_{j})}{h}.$ (27)
Having computed the quasi-normal frequencies, we can also obtain their corresponding amplitudes using the Prony fit. For this, we need to define a $(p\times N_{1})$ matrix $A_{ij}=z_{i-1}^{j}$ and a vector $v=[x_{1},x_{2},\ldots,x_{N_{1}}]$. We get the amplitudes by calculating $C=[C_{1},C_{2},\ldots,C_{p}]=(A^{\dagger}A)^{-1}A^{\dagger}v$. These amplitudes are generally complex numbers; however, we have reported only the absolute values in the appendix B.
In the method discussed above, we get $p$-number of solutions, but most of them are numerical artefacts and not actual QNMs of the system. To mitigate this issue, we can repeat the above steps with different values of $h$ and $N_{1}$ and accept only the results common to both sets of solutions (with appropriate tolerance) as the actual QNMs.
The time-domain and Prony fit method works well for most effective potentials which do not have any discontinuities. But this method often fails to return the $\omega_{QNM}$ with low relative amplitude (we notice this feature for odd overtones in Sec. III).
A.2 The spectral method
Using the corresponding equation for a perturbation (scalar or axial gravitational), another method known as the spectral method, may also be employed to find the QNMs of the system. In our analysis, we will use the Chebyshev grid and thus we have to map our coordinate to a finite interval, i.e. [1,1][-1,1]. The tortoise coordinate $l\in(-\infty,+\infty)$ may be mapped to $x\in[-1,1]$ using the transformation $x=\frac{2}{\pi}\arctan(l)$ Batic and Dutykh (2024).
$\frac{d^{2}\psi}{dl^{2}}+(\omega^{2}-V(l))\psi=0$ (28)
$\psi(x)=exp\left(i\omega\frac{2l}{\pi}\arctan(l)\right)\phi(x)$ (29)
$\phi(x)\approx\sum_{i=0}^{N}\phi(x_{i})C_{i}(x)$ (30)
with $x_{i}$ being the collocation points (Chebyshev roots) given by $x_{i}=\frac{(2i+1)\pi}{2(N+1)}$ (i=0,1,2,,N)(i={0,1,2,...,N}).
Thus, we can express the master equation in the form
$(M_{a})_{ij}=c_{0,a}D^{(0)}_{ij}+c_{1,a}D^{(1)}_{ij}+c_{2,a}D^{(2)}_{ij}\hskip 28.45274pta={1,2,...}$ (32)
$(M_{0}+\omega M_{1}+\omega^{2}M_{2}+...+\omega^{p}M_{p})\phi=0$ (33)
We can consider the example of scalar perturbation in the ultra-static Bronnikov–Ellis geometry (Eqn. (5)), where, after using the ansatz (Eqn. (29)) we have
$c_{0}=$ $-m(m+1)\left(\tan^{n}\left(\frac{\pi x}{2}\right)+1\right)^{-2/n}-\frac{(n-1)\tan^{n-2}\left(\frac{\pi x}{2}\right)}{\left(\tan^{n}\left(\frac{\pi x}{2}\right)+1\right)^{2}}-\left(x^{2}-1\right)\omega^{2}$
$+\frac{2i\pi\omega\cos(\pi x)-\omega\sin(\pi x)(2\pi x\omega+(\omega+i\pi)\sin(\pi x))}{\pi^{2}}+\frac{2i\omega}{\pi}$
$c_{1}=\frac{2i\cos^{2}\left(\frac{\pi x}{2}\right)(2\pi x\omega+(2\omega+i\pi)\sin(\pi x))}{\pi^{2}}\hskip 85.35826ptc_{2}=\frac{4\cos^{4}\left(\frac{\pi x}{2}\right)}{\pi^{2}}$
Now we can take the coefficient of $\omega^{0}$ in $c_{0}$ to get $c_{0,0}$, the coefficient of $\omega^{1}$ in $c_{0}$ to get $c_{1,0}$ and so on.
A.3 Other methods: merits and demerits
[…]
Appendix B QNMs for ultrastatic, generalized Bronnikov–Ellis wormhole
@@
In this ultra-static spacetime, if $(\omega_{R}-i\omega_{I})$ is a solution, then $(-\omega_{R}-i\omega_{I})$ is also a solution, both having the same amplitude. In the tables below, we show only the solution with a positive real part and double its amplitude (accounting for the amplitude of both solutions).
@@
0 0 0.681370.61775i0.68137-0.61775i $0.68402-0.61594i^{*}$ - - - -
1 0.496832.17988i0.49683-2.17988i - - - - -
1 0 1.572710.52970i1.57271-0.52970i $1.57274-0.52969i^{*}$ - - - -
1 1.259571.70681i1.25957-1.70681i - - - - -
2 0 2.546660.51267i2.54666-0.51267i 2.549250.51285i2.54925-0.51285i $0.35745915$ 1.737690.30514i1.73769-0.30514i $1.73771-0.30513i^{*}$ $0.2689666$
1 2.353001.57522i2.35300-1.57522i - - 1.712391.02010i1.71239-1.02010i - -
2 1.946412.75472i1.94641-2.75472i - - 1.524952.03050i1.52495-2.03050i - -
3 0 3.534290.50692i3.53429-0.50692i 3.534430.50686i3.53443-0.50686i $0.30122677$ 2.952410.40998i2.95241-0.40998i 2.952530.41006i2.95253-0.41006i $0.29811729$
1 3.390381.53568i3.39038-1.53568i - - 2.866431.25767i2.86643-1.25767i - -
2 3.099882.62324i3.09988-2.62324i 3.096822.62481i3.09682-2.62481i $0.16837004$ 2.660562.19085i2.66056-2.19085i 2.675182.17517i2.67518-2.17517i $0.171151$
4 0 4.527050.50433i4.52705-0.50433i 4.527340.50424i4.52734-0.50424i $0.26517902$ 4.076250.44911i4.07625-0.44911i 4.076460.44904i4.07646-0.44904i $0.2640608$
1 4.415361.52340i4.41536-1.52340i - - 3.994321.36285i3.99432-1.36285i - -
2 4.190232.57208i4.19023-2.57208i 4.190752.57242i4.19075-2.57242i $0.1397532$ 3.801002.30237i3.80100-2.30237i 3.798162.30549i3.79816-2.30549i $0.14061934$
5 0 5.522310.50296i5.52231-0.50296i 5.522820.50282i5.52282-0.50282i $0.23964086$ 5.154820.46725i5.15482-0.46725i 5.155240.46714i5.15524-0.46714i $0.239214$
1 5.429961.51789i5.42996-1.51789i 5.417961.52258i5.41796-1.52258i $2.009\times 10^{-5}$ 5.073061.40721i5.07306-1.40721i - -
2 5.247052.54683i5.24705-2.54683i 5.251232.54327i5.25123-2.54327i $0.12190640$ 4.907702.37002i4.90770-2.37002i 4.908152.37157i4.90815-2.37157i $0.12362226$
6 0 6.518970.50214i6.51897-0.50214i 6.519800.50195i6.51980-0.50195i $0.22033499$ 6.208820.47711i6.20882-0.47711i 6.209540.47695i6.20954-0.47695i $0.2201443$
1 6.442461.50964i6.44246-1.50964i 6.443131.53221i6.44313-1.53221i $1.886\times 10^{-5}$ 6.137291.43537i6.13729-1.43537i - -
2 6.286462.53346i6.28646-2.53346i 6.288192.53269i6.28819-2.53269i $0.11151401$ 5.991512.40927i5.99151-2.40927i 5.992672.40806i5.99267-2.40806i $0.1115693$
7 0 7.516490.50162i7.51649-0.50162i 7.519340.50106i7.51934-0.50106i $0.20454891$ 7.248180.48307i7.24818-0.48307i 7.249310.48285i7.24931-0.48285i $0.20497677$
1 7.449331.50895i7.44933-1.50895i 7.455761.50817i7.45576-1.50817i $2.309\times 10^{-5}$ 7.184381.45261i7.18438-1.45261i - -
2 7.315412.52544i7.31541-2.52544i 7.319852.52412i7.31985-2.52412i $0.10156691$ 7.056352.43256i7.05635-2.43256i 7.057852.43166i7.05785-2.43166i $0.10294975$
Table 8: $\omega_{\text{QNM}}$ values for generalised Bronnikov–Ellis wormhole ($n=2$) under scalar perturbations for $m=0$ - 77 and axial perturbations for $m=2$ - 77. [The QNM spectra for the $n=2$ case were published in Kim (2008); Batic and Dutykh (2024); Blázquez-Salcedo et al. (2018a); Alfaro et al. (2024)]